# Streaming Max-Cut in General Metrics

### v1

@@ 1.3 Related Work
Under streaming model, for graph streams, it is shown that breaking the $2$-approximation barrier requires $\Omega(n)$ space, even in insertion-only streams and when the edges are present in a random order [KK19]. In contrast, for geometric streams in $\mathbb{R}^{d}$, it is possible to obtain $(1+\epsilon)$-approximation in polylogarithmic space [CJK23]. Recently, this result has been extended to provide not only an approximate value but also an oracle to return an approximate $\operatorname{Max-Cut}$ solution [MW25]. Furthermore, Dong et al. [DPV25] recently gave a learning-augmented streaming algorithm that approximates the value of $\operatorname{Max-Cut}$ of a general graph with an approximation ratio slightly better than $2$.
@@ 2 Preliminaries
For $m\in\mathbb{N}$, write $m]:=\{1,2,\ldots,m\}$. Let $(V,\operatorname{dist})$ by the underlying metric space. Given a point set $P$, Given a point set $P$, the aspect ratio of $P$ is defined as the ratio between its largest and smallest non-zero pairwise distances. For some $\alpha\geq 1$, we say a non-negative real number $E$ is $\alpha$-approximate to $\operatorname{Max-Cut}(P)$ for dataset $P$, if $E\leq\operatorname{Max-Cut}(P)\leq(1+\epsilon)E$. We interpret a set as a multi-set throughout. The cut function $\operatorname{cut}:V\times V\to\mathbb{R}_{+}$ is defined as $\operatorname{cut}(S,T):=\sum_{x\in S}\sum_{y\in T}\operatorname{dist}(x,y)$ for $S,T\subseteq V$. For a subset $S\subseteq V$ and a dataset $P\subseteq V$, write $\operatorname{cut}_{P}(S):=\operatorname{cut}(S,P\setminus S)$. The $\operatorname{Max-Cut}$ value of a dataset $P\subseteq V$ is defined as
@@
3 An Algorithm for Insertion-Only Streams
In this section, we present an algorithm to approximate the value of $\operatorname{Max-Cut}$ for insertion-only streams and establish the following theorem.
@@ 3.1 Some Tools
We need the following lemma from [CJK23].
Lemma 3.2 ([CJK23]).
@@
We also require the following coreset tool for the $1$-median problem. Given a set $Q\subseteq V$ and a point $y\in V$, the $1$-median cost of $y$ with respect to $Q$ is defined as $\sum_{y^{\prime}\in Q}\operatorname{dist}(y^{\prime},y)$.
@@ 3.2 The Algorithm
We first describe our importance sampling algorithm. Given a data stream $p_{1},\dots,p_{n}$ of points in $P$, let $P_{i}:=\{p_{1},\ldots,p_{i}\}$ be the prefix of the first $i$ points. For each $p_{i}$, we define the prefix-based quantity $q^{\prime}(p_{i}):=\sum_{x\in P_{i}}\operatorname{dist}(p_{i},x)$. We maintain an $\epsilon$-coreset for the $1$-median objective $\operatorname{med}(P_{i},x)$, and use it to maintain a single sample $s\in P$. Specifically, when $p_{i}$ arrives, we replace the current sample with $p_{i}$ with probability proportional to $\frac{q^{\prime}(p_{i})}{\sum_{j\leq i}q^{\prime}(p_{j})}$, and otherwise keep the current sample. The full algorithm is given in Algorithm 1.
@@
2 initialize $(C,w_{C}^{\prime})$ as an $\epsilon$-coreset for 11-median (initially $C=\emptyset$, $w_{C}^{\prime}(c)=0$ for any $c$)
@@ 9    $\beta_{t}\leftarrow\frac{1}{K}\cdot\frac{\hat{R}_{t}}{\hat{Q}_{t}}$
Given the above sampling procedure, we can now describe our algorithm for estimating $\operatorname{Max-Cut}(P)$.
@@
4Invoke Algorithm 1 in parallel over the stream to obtain a weighted sample set $(S,w_{S})$ of size $m$, where $m\geq\Theta(\epsilon^{-4}K^{8})$, and $w_{S}(s)=w$, when $(s,w)$ is returned by Algorithm 1.
5 return $\eta:=\frac{\mathrm{Max\text{-}Cut}(S)}{m^{2}}$
@@
3.3 The Analysis
Now we analyze Algorithm 1. Define $R_{t}:=\sum_{i=1}^{t}\operatorname{dist}(p_{t},p_{i})$ and $Q_{t}:=\sum_{i=1}^{t}\sum_{j=1}^{t}\operatorname{dist}(p_{i},p_{j})$. Without less of generality, we assume $Q_{n}>0$.
@@ 3.3 The Analysis
$(1-\epsilon)\cdot Q_{t}$ $\leq\hat{Q}_{t}\leq(1+\epsilon)\cdot Q_{t}.$ (3)
@@
Note that $R_{t}=\sum_{i=1}^{t}\operatorname{dist}(p_{t},p_{i})$ and $\hat{R}_{t}=\sum_{c\in C}\operatorname{dist}(c,p_{t})w_{C}^{\prime}(c)$. Thus Equation 2 holds for all $t$.
Furthermore, since $\hat{Q}_{t}=\sum_{i=1}^{t}\hat{R}_{i}$, from Lemma 3.5, we get
$\hat{Q}_{t}=\sum_{i=1}^{t}\hat{R}_{i}\geq(1-\epsilon)\cdot\sum_{i=1}^{t}R_{t}=(1-\epsilon)\cdot Q_{t}$
Similarly, we have that $\hat{Q}_{t}\leq(1+\epsilon)\cdot Q_{t}$. Thus Equation 3 holds for all $t$. ∎
In the following two lemmas, we give upper bounds for the quantity $\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$.
@@
Thus, we have that
$\frac{1}{t}\geq\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}.$
This finishes the proof of the lemma. ∎
Lemma 3.7.
For any $t$, if $Q_{t}>0$, then the following inequality holds:
Proof.
We first note that it suffices to prove that for every $a\in\{t+1,\dots,n\}$,
$\frac{\operatorname{dist}(p_{t},p_{a})}{2\sum_{j=1}^{t}\operatorname{dist}(p_{a},p_{j})}\leq\frac{2R_{t}}{Q_{t}}.$
Indeed, if this holds, then
$\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$
$\leq$ $\frac{R_{t}+\sum_{i=t+1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{t}+2\sum_{i=t+1}^{n}\sum_{j=1}^{t}\operatorname{dist}(p_{i},p_{j})}$
$\leq$ $\frac{2R_{t}}{Q_{t}},$
We will show that this leads to a contradiction.
Using the definition of $Q_{t}$, we have:
Combining this with Equation 4, we can deduce:
$\sum_{j=1}^{t}\operatorname{dist}(p_{a},p_{j})<\frac{1}{2}t\cdot\operatorname{dist}(p_{t},p_{a}).$
Let us analyze $\frac{R_{t}}{Q_{t}}$ in this scenario. We have:
$\frac{R_{t}}{Q_{t}}$ $\geq\frac{\sum_{i=1}^{t}\left(\operatorname{dist}(p_{t},p_{a})-\operatorname{dist}(p_{i},p_{a})\right)}{\sum_{i=1}^{t}\sum_{j=1}^{t}\left(\operatorname{dist}(p_{i},p_{a})+\operatorname{dist}(p_{j},p_{a})\right)}$
$=\frac{t\cdot\operatorname{dist}(p_{t},p_{a})-\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}{2t\cdot\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}$
$>\frac{t\cdot\operatorname{dist}(p_{t},p_{a})-\frac{1}{2}t\cdot\operatorname{dist}(p_{t},p_{a})}{2t\cdot\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}$
$=\frac{\operatorname{dist}(p_{t},p_{a})}{4\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}.$
However, this contradicts Equation 4. Therefore, no such $a$ satisfying Equation 4 can exist, and we conclude:
$\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}\leq\frac{2R_{t}}{Q_{t}}.$
This completes the proof. ∎
Recall that $s$ is the sample returned by Algorithm 1. We show that for any point $x\in P$, the probability that $s=x$ is at least $\Omega\left(\frac{1}{K}\cdot\tfrac{q(x)}{Q}\right)$, where $q(x)=\sum_{y\in P}\operatorname{dist}(x,y)$ and $Q=Q_{n}$. This bound allows us to invoke Lemma 3.2 and thereby ensure the correctness of our algorithm. Formally, we establish the following lemma.
Lemma 3.8.
Suppose that the inequalities in Lemma 3.5 holds for all $t$. For any $p_{t}\in P$, if Algorithm 1 returns $(p_{t},w)$, the following inequality holds:
Proof.
If $s=p_{t}$, then $p_{t}$ is sampled when it comes, and all subsequent points are not sampled, so the sampling probability $\Pr[s=p_{t}]$ is given by:
$\Pr[s=p_{t}]=\beta_{t}\cdot\prod_{i=t+1}^{n}(1-\beta_{i}),$
which is the value of $w$. In the following, we consider the case that $Q_{t}>0$ first. Using the inequality $\prod_{i}(1-x_{i})\geq 1-\sum_{i}x_{i}$ for $0\leq x_{i}\leq 1$, we have:
$\Pr[s=p_{t}]\geq\beta_{t}\cdot\left(1-\sum_{i=t+1}^{n}\beta_{i}\right).$
Substituting the definition of $\beta_{i}$, this becomes:
$\Pr[s=p_{t}]\geq\beta_{t}\cdot\left(1-\frac{1}{K}\cdot\sum_{i=t+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}\right).$
Applying Lemma 3.5, which bounds $\frac{\hat{R}_{t}}{\hat{Q}_{t}}$ in terms of $\frac{R_{t}}{Q_{t}}$, we get:
$\beta_{t}=\frac{1}{K}\cdot\frac{\hat{R}_{t}}{\hat{Q}_{t}}\geq\frac{1}{K}\cdot\frac{1-\epsilon}{1+\epsilon}\cdot\frac{R_{t}}{Q_{t}}.$
Using Lemma 3.7, which bounds $\frac{2R_{t}}{Q_{t}}$ in terms of $\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$, we further obtain:
$\beta_{t}\geq\frac{1}{2K}\cdot\frac{1-\epsilon}{1+\epsilon}\cdot\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}.$
To bound the sum $\sum_{i=t+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}$, note that:
$\frac{\hat{R}_{i}}{\hat{Q}_{i}}\leq\frac{1+\epsilon}{1-\epsilon}\cdot\frac{R_{i}}{Q_{i}},$
and as $Q_{k}=\sum_{i=1}^{k}\sum_{j=1}^{k}\operatorname{dist}(p_{i},p_{j})=\sum_{i=1}^{k-1}\sum_{j=1}^{k-1}\operatorname{dist}(p_{i},p_{j})+\sum_{j=1}^{k}\operatorname{dist}(p_{k},p_{j})+\sum_{i=1}^{k-1}\operatorname{dist}(p_{i},p_{k})=Q_{k-1}+2R_{k}$, the following holds:
Using the fact that $(1-x)\leq e^{-x}$, we derive:
Since $1\leq\operatorname{dist}(p_{i},p_{j})\leq\Delta$ for all $i,j$, it follows that:
$\frac{2}{\Delta n^{2}}\leq\frac{Q_{2}}{Q_{n}}.$
Combining the above, we have $\frac{2}{\Delta n^{2}}\leq e^{-2\cdot\sum_{k=3}^{n}\frac{R_{k}}{Q_{k}}}$ and thus
$\sum_{k=3}^{n}\frac{R_{k}}{Q_{k}}\leq\ln(\Delta n^{2}).$
Furthermore,
$\sum_{k=2}^{n}\frac{\hat{R}_{k}}{\hat{Q}_{k}}\leq 1+\frac{1+\epsilon}{1-\epsilon}\cdot\ln(\Delta n^{2})\leq\frac{K}{2}.$
Substituting this back, we obtain:
Simplifying, we find:
$\Pr[s=p_{t}]\geq\frac{1-\epsilon}{4K(1+\epsilon)}\cdot\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}.$
Now, we consider the case that $Q_{t}=0$. Let $j\geq t$ be the largest index that $Q_{j}=0$. Then for any $a,b\leq j$, $\operatorname{dist}(p_{a},p_{b})=0$. We have
$\Pr[s=p_{t}]$ $=\beta_{t}\cdot\prod_{i=t+1}^{n}(1-\beta_{i})$
$=\frac{1}{t}\cdot\prod_{i=t+1}^{j}(1-\frac{1}{i})\cdot\prod_{i=j+1}^{n}(1-\beta_{i})$
$\geq\frac{1}{j}\cdot\left(1-\frac{1}{K}\cdot\sum_{i=j+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}\right).$
Similarly as above, we get
$\sum_{i=j+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}\leq\frac{K}{2}.$
Combining with Lemma 3.6, we get
@@
Now we are ready to finish the proof of Theorem 3.1.
Proof of Theorem 3.1.
Set $\lambda=\tfrac{4K(1+\epsilon)}{1-\epsilon}$; note that $K=O(\lambda)$. Our algorithm invokes Algorithm 1 $m$ times with $m=O(\epsilon^{-4}K^{8})=O(\epsilon^{-4}\lambda^{8})$ to obtain the sample set $S$ and corresponding weights. By Lemma 3.5, for each invocation of Algorithm 1, with probability at least $1-\tfrac{1}{n^{2}}$, the inequalities in Equation 2 hold for all $t\leq n$. Under this condition, Lemma 3.8 ensures that each sample $s$ with weight $w$ from Algorithm 1 satisfies the precondition of Lemma 3.2, where the distribution $\mathcal{D}$ is given by $p_{x}=\Pr[s=x]\geq\tfrac{1}{\lambda}\cdot\tfrac{q(x)}{Q}$, with $q(x):=\sum_{y\in P}\operatorname{dist}(x,y)$ and $Q:=\sum_{x\in P}q(x)$. Furthermore, the weight is $w_{S}(x)=\Pr[s=x]$. Since we invoke Algorithm 1 for $m\ll n$ times, Lemma 3.2 implies that with probability at least $1-\tfrac{m}{n^{2}}-0.1>\tfrac{2}{3}$, the estimate $\eta$ is a $(1+\epsilon)$-approximation of $\operatorname{Max-Cut}(P)$.
To bound the space complexity, note that by Lemma 3.4, a single invocation of Algorithm 1 requires $\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$ space to maintain the coreset. Since we invoke Algorithm 1 $m$ times, the total space for all coresets is still $m\cdot\operatorname{poly}(\epsilon^{-1}\log(n\Delta))=\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$. After obtaining the set $S$, we query all $O(m^{2})$ pairwise distances to estimate the weighted max cut, which also requires $O(m^{2})=\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$ space. Hence, the overall space complexity is $\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$. ∎
4 An Algorithm for Sliding-Window Streams

### v2

@@ 1.3 Related Work
Under streaming model, for graph streams, it is shown that breaking the $2$-approximation barrier requires $\Omega(n)$ space, even in insertion-only streams and when the edges are present in a random order [KK19]. In contrast, for geometric streams in $\mathbb{R}^{d}$, it is possible to obtain $(1+\epsilon)$-approximation in polylogarithmic space [CJK23]. Recently, this result has been extended to provide not only an approximate value but also an oracle to return an approximate $\operatorname{Max-Cut}$ solution [MW26]. Furthermore, Dong et al. [DPV25] recently gave a learning-augmented streaming algorithm that approximates the value of $\operatorname{Max-Cut}$ of a general graph with an approximation ratio slightly better than $2$.
@@ 2 Preliminaries
For $m\in\mathbb{N}$, write $m]:=\{1,2,\ldots,m\}$. Let $(V,\operatorname{dist})$ by the underlying metric space. Given a point set $P$, Given a point set $P$, the aspect ratio of $P$ is defined as the ratio between its largest and smallest non-zero pairwise distances. For some $\alpha\geq 1$, we say a non-negative real number $E$ is $\alpha$-approximate to $\operatorname{Max-Cut}(P)$ for dataset $P$, if $E\leq\operatorname{Max-Cut}(P)\leq\alpha E$. We interpret a set as a multi-set throughout. The cut function $\operatorname{cut}:V\times V\to\mathbb{R}_{+}$ is defined as $\operatorname{cut}(S,T):=\sum_{x\in S}\sum_{y\in T}\operatorname{dist}(x,y)$ for $S,T\subseteq V$. For a subset $S\subseteq V$ and a dataset $P\subseteq V$, write $\operatorname{cut}_{P}(S):=\operatorname{cut}(S,P\setminus S)$. The $\operatorname{Max-Cut}$ value of a dataset $P\subseteq V$ is defined as
@@
3 Algorithms for Insertion-Only Streams
In this section, we present an algorithm to approximate the value of $\operatorname{Max-Cut}$ for insertion-only streams and establish the following theorem. The sliding-window result is based on this insertion-only one, and will be presented in Section 4.
@@ 3.1 Some Tools
Our high-level framework is the following importance sampling suggested by [CJK23]. Roughly speaking, one samples each data point $x$ with probability (approximately) proportional to the sum of distances from $x$ to every other points, then around $\operatorname{poly}\log n$ samples is enough to approximate the $\operatorname{Max-Cut}$ value.
Lemma 3.2 (Importance Sampling [CJK23]).
@@
We also require the following coreset tool for $1$-median clustering. Given a set $U\subseteq V$ and a point $y\in V$, the $1$-median cost of $y$ with respect to $U$ is defined as $\sum_{y^{\prime}\in U}\operatorname{dist}(y^{\prime},y)$.
@@ 3.2 The Algorithm
We propose a variant of importance sampling as in Lemma 3.2. Our importance sampling uses some $q^{\prime}$ as the importance score, which is defined with respect to only a prefix of the stream, instead of $q$. Specifically, given a data stream $p_{1},\dots,p_{n}$ of points in $P$, let $P_{i}:=\{p_{1},\ldots,p_{i}\}$ be the prefix of the first $i$ points. For each $p_{i}$, we define $q^{\prime}(p_{i}):=\sum_{x\in P_{i}}\operatorname{dist}(p_{i},x)$ with respect to the prefix. We maintain an $\epsilon$-coreset for the $1$-median objective $\operatorname{med}(P_{i},x)$ (as in Lemma 3.4), and use it to maintain a single sample $s\in P$: when $p_{i}$ arrives, we replace the current sample with $p_{i}$ with probability (approximately) proportional to $\frac{q^{\prime}(p_{i})}{\sum_{j\leq i}q^{\prime}(p_{j})}$, and otherwise keep the current sample. We give the full algorithm in Algorithm 1.
@@
2 initialize $(C,w_{C}^{\prime})$ as an $\epsilon$-coreset for $1$-median (initially $C=\emptyset$, $w_{C}^{\prime}(c)=0$ for any $c$)
@@ 9    $\beta_{t}\leftarrow\frac{1}{K}\cdot\frac{\hat{R}_{t}}{\hat{Q}_{t}}$
The final algorithm for estimating $\operatorname{Max-Cut}(P)$ is given in Algorithm 2, where we plug in Lemma 3.2 the approximate estimate as in Algorithm 1.
@@
4let $K$ be as in Algorithm 1, and let $\lambda:=\tfrac{4K(1+\epsilon)}{1-\epsilon}$
5
6maintain $m:=\Theta(\epsilon^{-4}K^{8})$ independent runs of Algorithm 1 during the update of the stream
7
8query each run to obtain a sample $(s_{i},w_{i})$ to define $(S,w_{S}):=(\{s_{i}\}_{i=1}^{m},\{w_{i}\}_{i=1}^{m})$
9 return $\eta:=\frac{\mathrm{Max\text{-}Cut}(S)}{m^{2}}$ as in Lemma 3.2
@@
3.3 Proof of Theorem 3.1
We start with the analysis of Algorithm 1. Define $R_{t}:=\sum_{i=1}^{t}\operatorname{dist}(p_{t},p_{i})$ as the sum of distances from $p_{t}$ to the prefix $P_{t}$, and $Q_{t}:=\sum_{i=1}^{t}\sum_{j=1}^{t}\operatorname{dist}(p_{i},p_{j})$ be the total sum of all distances among the prefix $P_{t}$. Without less of generality, we assume $Q_{n}>0$. Recall that $s$ is the sample returned by Algorithm 1. We show that for any point $x\in P$, the probability that $s=x$ is at least $\Omega\left(\frac{1}{K}\cdot\tfrac{q(x)}{Q}\right)$, where $q(x)=\sum_{y\in P}\operatorname{dist}(x,y)$ and $Q=Q_{n}$. This bound essentially says that the importance score $q^{\prime}(x)$ that is defined only with respect to the prefix, is a good approximation for the $q(x)$’s defined with respect to the full stream. This further allows us to invoke Lemma 3.2 and thereby ensure the correctness of our algorithm. Formally, we establish the following key lemma.
Lemma 3.5.
Let $(s,w)$ denote the return value of Algorithm 1. With probability $1-1/n^{2}$, for every $p_{t}\in P$, the following inequality holds:
The proof of this lemma is postponed to Section 3.4, and we first conclude the proof of Theorem 3.1 assuming Lemma 3.5 is correct.
Proof of Theorem 3.1.
Note that $K=O(\lambda)$. Our algorithm invokes Algorithm 1 $m$ times with $m=O(\epsilon^{-4}K^{8})=O(\epsilon^{-4}\lambda^{8})$ to obtain the sample set $S$ and corresponding weights. Lemma 3.5 ensures that each sample $s$ with weight $w$ from Algorithm 1 satisfies the premises of Lemma 3.2, where the distribution $\mathcal{D}$ is given by $p_{x}=\Pr[s=x]\geq\tfrac{1}{\lambda}\cdot\tfrac{q(x)}{Q}$, with $q(x):=\sum_{y\in P}\operatorname{dist}(x,y)$ and $Q:=\sum_{x\in P}q(x)$. Furthermore, the weight is $w_{S}(x)=\Pr[s=x]$. Since we invoke Algorithm 1 for $m\ll n$ times, Lemma 3.2 implies that with probability at least $1-\tfrac{m}{n^{2}}-0.1>\tfrac{2}{3}$, the estimate $\eta$ is a $(1+\epsilon)$-approximation of $\operatorname{Max-Cut}(P)$.
To bound the space complexity, note that by Lemma 3.4, a single invocation of Algorithm 1 requires $\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$ space to maintain the coreset. Since we invoke Algorithm 1 $m$ times, the total space for all coresets is $m\cdot\operatorname{poly}(\epsilon^{-1}\log(n\Delta))=\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$. After obtaining the set $S$, we query all $O(m^{2})$ pairwise distances to estimate the weighted max cut, which requires $O(m^{2})=\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$ space. Hence, the overall space complexity is $\operatorname{poly}(\epsilon^{-1}\log(n\Delta))$. ∎
3.4 Proof of Key Lemma 3.5
Proof.
Since we consider $s=p_{t}$, then $p_{t}$ is sampled when it comes, and all subsequent points are not sampled, so the sampling probability $\Pr[s=p_{t}]$ is given by: $w=\Pr[s=p_{t}]=\beta_{t}\cdot\prod_{i=t+1}^{n}(1-\beta_{i})$.
Case I: $Q_{t}>0$.
In the following, we consider the case that $Q_{t}>0$. Using the inequality $\prod_{i}(1-x_{i})\geq 1-\sum_{i}x_{i}$ for $0\leq x_{i}\leq 1$, we have: $\Pr[s=p_{t}]\geq\beta_{t}\cdot\left(1-\sum_{i=t+1}^{n}\beta_{i}\right)$. Substituting the definition of $\beta_{i}$, this becomes:
$\Pr[s=p_{t}]\geq\beta_{t}\cdot\left(1-\frac{1}{K}\cdot\sum_{i=t+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}\right).$ (2)
Next, we bound $\frac{\hat{R}_{i}}{\hat{Q}_{i}}$ in the following lemma.
@@
$(1-\epsilon)Q_{t}$ $\leq\hat{Q}_{t}\leq(1+\epsilon)Q_{t}.$ (4)
@@
Note that $R_{t}=\sum_{i=1}^{t}\operatorname{dist}(p_{t},p_{i})$ and $\hat{R}_{t}=\sum_{c\in C}\operatorname{dist}(c,p_{t})w_{C}^{\prime}(c)$. Thus (3) holds for all $t$. Furthermore, since $\hat{Q}_{t}=\sum_{i=1}^{t}\hat{R}_{i}$, applying (3) we get $\hat{Q}_{t}=\sum_{i=1}^{t}\hat{R}_{i}\geq(1-\epsilon)\cdot\sum_{i=1}^{t}R_{t}=(1-\epsilon)Q_{t}$. Similarly, $\hat{Q}_{t}\leq(1+\epsilon)Q_{t}$. Hence, (4) holds for all $t$. This finishes the proof of Lemma 3.6. ∎
Continue with (2), applying Lemma 3.6 which bounds $\frac{\hat{R}_{t}}{\hat{Q}_{t}}$ in terms of $\frac{R_{t}}{Q_{t}}$, we get
$\beta_{t}=\frac{1}{K}\cdot\frac{\hat{R}_{t}}{\hat{Q}_{t}}\geq\frac{1}{K}\cdot\frac{1-\epsilon}{1+\epsilon}\cdot\frac{R_{t}}{Q_{t}}.$ (5)
Next, we establish the following Lemma 3.7, to bound $\frac{R_{t}}{Q_{t}}$ with respect to $\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$.
Lemma 3.7.
For any $t$, if $Q_{t}>0$, then the following inequality holds:
Proof.
We first note that it suffices to prove that for every $a\in\{t+1,\dots,n\}$, it holds that $\frac{\operatorname{dist}(p_{t},p_{a})}{2\sum_{j=1}^{t}\operatorname{dist}(p_{a},p_{j})}\leq\frac{2R_{t}}{Q_{t}}$. Indeed, if this holds, then
$\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$
$\leq$ $\frac{R_{t}+\sum_{i=t+1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{t}+2\sum_{i=t+1}^{n}\sum_{j=1}^{t}\operatorname{dist}(p_{i},p_{j})}\leq\frac{2R_{t}}{Q_{t}},$
We will show that this leads to a contradiction. Using the definition of $Q_{t}$, we have:
Combining this with Equation 6, we can deduce: $\sum_{j=1}^{t}\operatorname{dist}(p_{a},p_{j})<\frac{1}{2}t\cdot\operatorname{dist}(p_{t},p_{a})$. Let us analyze $\frac{R_{t}}{Q_{t}}$ in this scenario. We have:
$\frac{R_{t}}{Q_{t}}$ $\geq\frac{\sum_{i=1}^{t}\left(\operatorname{dist}(p_{t},p_{a})-\operatorname{dist}(p_{i},p_{a})\right)}{\sum_{i=1}^{t}\sum_{j=1}^{t}\left(\operatorname{dist}(p_{i},p_{a})+\operatorname{dist}(p_{j},p_{a})\right)}=\frac{t\cdot\operatorname{dist}(p_{t},p_{a})-\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}{2t\cdot\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}$
$>\frac{t\cdot\operatorname{dist}(p_{t},p_{a})-\frac{1}{2}t\cdot\operatorname{dist}(p_{t},p_{a})}{2t\cdot\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}=\frac{\operatorname{dist}(p_{t},p_{a})}{4\sum_{i=1}^{t}\operatorname{dist}(p_{i},p_{a})}.$
However, this contradicts Equation 6. Therefore, no such $a$ satisfying Equation 6 can exist, and we conclude: $\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}\leq\frac{2R_{t}}{Q_{t}}$. This completes the proof of Lemma 3.7. ∎
Continue with (5), using Lemma 3.7 which bounds $\frac{2R_{t}}{Q_{t}}$ in terms of $\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$, we further obtain:
$\beta_{t}\geq\frac{1}{2K}\cdot\frac{1-\epsilon}{1+\epsilon}\cdot\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}.$ (7)
We still need to bound the sum $\sum_{i=t+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}$ in order to bound (2). Note that $\frac{\hat{R}_{i}}{\hat{Q}_{i}}\leq\frac{1+\epsilon}{1-\epsilon}\cdot\frac{R_{i}}{Q_{i}}$, and as $Q_{k}=\sum_{i=1}^{k}\sum_{j=1}^{k}\operatorname{dist}(p_{i},p_{j})=\sum_{i=1}^{k-1}\sum_{j=1}^{k-1}\operatorname{dist}(p_{i},p_{j})+\sum_{j=1}^{k}\operatorname{dist}(p_{k},p_{j})+\sum_{i=1}^{k-1}\operatorname{dist}(p_{i},p_{k})=Q_{k-1}+2R_{k}$, the following holds:
Using the fact that $(1-x)\leq e^{-x}$, we derive:
Since $1\leq\operatorname{dist}(p_{i},p_{j})\leq\Delta$ for all $i,j$, it follows that $\frac{2}{\Delta n^{2}}\leq\frac{Q_{2}}{Q_{n}}$. Combining the above, we have $\frac{2}{\Delta n^{2}}\leq e^{-2\cdot\sum_{k=3}^{n}\frac{R_{k}}{Q_{k}}}$ and thus $\sum_{k=3}^{n}\frac{R_{k}}{Q_{k}}\leq\ln(\Delta n^{2})$. Furthermore,
$\sum_{k=2}^{n}\frac{\hat{R}_{k}}{\hat{Q}_{k}}\leq 1+\frac{1+\epsilon}{1-\epsilon}\cdot\ln(\Delta n^{2})\leq\frac{K}{2}.$ (8)
Substituting this and (7) back to (2), we obtain:
This yields $\Pr[s=p_{t}]\geq\frac{1-\epsilon}{4K(1+\epsilon)}\cdot\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$, which concludes the $Q_{t}>0$ case.
Case II: $Q_{t}=0$.
We turn to the remaining case $Q_{t}=0$. Let $j\geq t$ be the largest index that $Q_{j}=0$. Then for any $a,b\leq j$, $\operatorname{dist}(p_{a},p_{b})=0$. We have
$\Pr[s=p_{t}]$ $=\beta_{t}\cdot\prod_{i=t+1}^{n}(1-\beta_{i})=\frac{1}{t}\cdot\prod_{i=t+1}^{j}(1-\frac{1}{i})\cdot\prod_{i=j+1}^{n}(1-\beta_{i})$
$\geq\frac{1}{j}\cdot\left(1-\frac{1}{K}\cdot\sum_{i=j+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}\right).$ (9)
Since $Q_{i}\neq 0$ for $i\geq j+1$, we can apply (8) to get
$\sum_{i=j+1}^{n}\frac{\hat{R}_{i}}{\hat{Q}_{i}}\leq\frac{K}{2}.$ (10)
To also lower bound $1/j$, we establish the following Lemma 3.8.
@@
Thus, we have that $\frac{1}{t}\geq\frac{\sum_{i=1}^{n}\operatorname{dist}(p_{t},p_{i})}{Q_{n}}$. This finishes the proof of Lemma 3.8. ∎
Finally, Substituting Lemma 3.8 and (10) to (9), we have
@@
This finishes the proof of the key Lemma 3.5. ∎
4 Algorithms for Sliding-Window Streams